Recent Log Entries

COMPLETED Install Hall C Turbo LCW Cooling Lines

Log Number: 3814795Logbook(s): Body: 

LCW hoses run and flow established. I would recommend
revisiting this work during SAD and installing a more robust
connection to the turbo. Presently it is a bard fitting and hose
clamps. Also installing isolation valves to all of the LCW taps
on the mezzanine.

COMMENT ON Crate Building

Log Number: 3814791Logbook(s): Body: 

Transform shipping crates requested by Rick Nelson, charge code
is RSRRF. Transformers presently located in NAB will need to be
moved to Bldg 36.

Back phasing bunch length measurement data

Log Number: 3814790Logbook(s): Entry Makers: kazimiBody: 

The bunch length measurement by back phasing in 4D00 spectrometer was done for Lasers A, B and C beams.
A save was done at the start of the work 21430 and another at the end 21433.
Fig 1: Initial rf setup
Fig 2,3: -90,+90 laser A
Fig 4,5: -91,+91 laser A
Fig 6: -90,+90 Harp laser A
Fig 7: -91, +91 Harp laser A
Fig 8: -90,-90 laser A
Fig 9: +90,+90 laser A
Fig 10: -90,-90 laser A Harp
Fig 11: +90,+90 laser A Harp
Fig 12: +90,+90 laser B Harp
Fig 13: +90,+90 laser C Harp
Fig 14: -90,-90 laser B Harp
Fig 15: -90,-90 laser C Harp
Fig 16: -90,-90 laser B
Fig 17: -90,-90 laser C

I will do the anal

Degaussing 4D00 magnet

Log Number: 3814789Logbook(s): Entry Makers: kazimiBody: 

I notice that correctors around 4D spectrometer have very high horizontal values. Degaussing the spectrometer magnet (Figure 1) and re-steering the line (Figure 2) lowered the corrector settings (Figure 3).
I know that our spectrometer magnet cycling routine for "Straight Ahead" needs some special dance to go the -4.8 Amps but I can not remember it exactly. I will email this to Gary Croke and Alicia. I would be good to fix the problem so that it works automatically.
I have left the magnet setting at setpoint #2 after degaussing (not the normal setpoint #1).

Autolog save

Log Number: 3814783Logbook(s): Tags: Body: saver ran from opsl632

Saved: kazimi 07-14-20 23:58:55 {A:1 B:5 C:5 E:1050 A=BS B=BS C=BS D=BS 0R08=0.00 0R08 dump in.

Autolog save

Log Number: 3814782Logbook(s): Tags: Body: saver ran from opsl632

Saved: kazimi 07-14-20 23:58:24 {A:1 B:5 C:5 E:1050 A=Tune B=BS C=BS D=BS 0R08=0.00 0R08 dump in.

CREW CHIEF SIGN IN

Log Number: 3814780Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 07-14-20
Time : 23:48:29
Day : Tuesday
Shift : Owl

Crew Chief : Frierson
Program Deputy: McGuckin
Team : D
Arm : RADCON
LERF Operator:
Operators :

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

MQD0R09 -- No Trim Card in Slot!

Log Number: 3814779Logbook(s): Entry Makers: moserBody: 

Linked log explains how MQD0R09 was not successfully cycling. Appeared calibration of card was off. Went to swap the card with a ready spare. Pulled the old trim card out and attempted to install a new one, but after over a dozen attempts could not get the card to properly seat with the backplane. The top rail appeared to be loose and wobbly about halfway into the rack.

Tried the original card again. First shot, went right in. Pulled it back out and compared the two --no obvious geometric differences. Tried the new card again, no luck.

COMPLETED ^LERF PSS Certification

Log Number: 3814778Logbook(s): Body: 

Completed Gun and Laser Chassis Functional Tests.

RF201, 202, 203 and Buncher were not tested.
This equipment will remain locked up until troubleshooting and repairs are complete and RF is needed.

W501B11 Negative Bulk Power Supply Tweaked

Log Number: 3814777Logbook(s): Entry Makers: moserBody: 

MQA{4, 6, 8}A21, off of W501B11 were failing to successfully cycle due to not making it to HMIN. Negative bulk power supply was outputting 28.4V, adjusted to 30.0. Quads cycling fine now.

MQA3A29 Trim Card Swapped

Log Number: 3814776Logbook(s): Entry Makers: moserBody: 

Magnet setpoint was stuck. Not taking demand, not cycling. Swapped trim card, all is well.

1L25-8 Vacuum Degredation at ODVH=20.5Mv/m

Log Number: 3814775Logbook(s): Entry Makers: kernsBody: 

Crew chief and I played with trying to keep this zone from dropping HV . Last week I found that cavity 8 would drop all eight out at anything above 10Mv/m. Gradually lowered from 20.5 Mv/m until gradient stopped affecting the vacuum with all the others at ODVH. Didn't change ODVH due to forgetfullnes and Lem put it back at 20.5 and it dropped out several dozen times until MCC made not of it. D. Moser will see how it holds up at 10Mv/m and we'll take it in the morning !!!

CEBAF Magnets from 3/24 Downloaded, Cycled

Log Number: 3814773Logbook(s): Entry Makers: moserBody: 

Used all save 21402 per the COVID restart plan, ATLis 20836.

First, selected "All Magnets"; grep -v 0[LIR]; grep BDL; Get Current Values. Download.... Performed a back-up all-save (in addition to auto-save before download). #21431

Next, cycled all box supplies. Then trims.

HYST AREA 6-2-1 Cycles Magnets

Log Number: 3814771Logbook(s): Body: 
HYST AREA Version 6-2-1 has FAILED to cycle:

MQD1L05 -- NL01B02 -- nl6mag -- Failed to set after 2 cycle attempts

MQA3A29 -- SA01B06 -- ea6mag -- Failed to set after 2 cycle attempts

MBDAS10H -- W102B10 -- wa1mag -- Failed to set after 2 cycle attempts

MQA4A12 -- W301B13 -- wa3mag -- Failed to set after 2 cycle attempts

MQA4A21 -- W501B11 -- wa5mag -- Failed to set after 2 cycle attempts
MQA6A21 -- W501B11 -- wa5mag -- Failed to set after 2 cycle attempts

Box Supplies Were Cycled

Log Number: 3814769Logbook(s): Tags: Body: 

The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
ARC1A Ramp Rate = 8.297 A/s
ARC2A Ramp Rate = 6.895 A/s
ARC3A Ramp Rate = 9.835 A/s
ARC4A Ramp Rate = 11.047 A/s
ARC5A Ramp Rate = 10.161 A/s
ARC6A Ramp Rate = 9.135 A/s
ARC7A Ramp Rate = 10.203 A/s
ARC8A Ramp Rate = 11.950 A/s
ARC9A Ramp Rate = 8.959 A/s
ARCAA Ramp Rate = 12.244 A/s
LAM1C Ramp Rate = 10.500 A/s
LAM3C Ramp Rate = 11.854 A/s
MYA2T01 Ramp Rate N/A
MYA4T01 Ramp Rate N/A
MYA6T01 Ramp Rate N/A
MYAAT01 Ramp Rate N/A
RSEP8A Ramp Rate = 10.471 A/s
RSEP9A Ramp Rate = 9.366 A/s

Pages